OpenStudy (anonymous):

Find all values of t in [0, 2π] for which |csct|=2

OpenStudy (anonymous):

csc = 2 if csc >= 0 and csc = -2 if csc < 0 csc = 1/sin so sin = 1/2 or -1/2 that give us pi/6 or 5pi/6

OpenStudy (anonymous):

Can you help with a few more?

OpenStudy (anonymous):

There are more answers those two are valid for sin = 1/2 and we have also 7pi/6 and 11pi/6

OpenStudy (anonymous):

I can try :D

OpenStudy (anonymous):

Thank you! :) Find all values of t in [0, 2π] for which |tan t|=1

OpenStudy (anonymous):